# Efficient Transformed Gaussian Process State-Space Models for Non-Stationary High-Dimensional Dynamical Systems

### v2

@@ III-B Approximate Bayesian Inference
To infer the latent variables $\{{\mathbf{x}}_{0:T},\mathbf{w}\}$ and the GP $\tilde{f}$, we employ the following variational distribution:
$q({\mathbf{x}}_{0:T},\tilde{f},\mathbf{w})\!=\!q(\tilde{f}),q(\mathbf{w}),q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1}),$ (21)
which mirrors the generative structure of the model in Eq. (18) and assumes independence between $\tilde{f}$ and $\mathbf{w}$. Here the variational distributions of ${\mathbf{x}}_{0}$ and $\mathbf{w}$ are defined as follows:
[…]
For $q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})$ in Eq. (21), rather than assuming a parameterized form—which would introduce additional variational parameters and hence increase optimization complexity [31]—we draw inspiration from recent works [37, 50] and incorporate the EnKF into the variational learning framework to infer the latent states (see the next subsection for details). Specifically, without explicitly defining the variational distribution, we adopt the following approximation:
Assumption 1.
$q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\approx p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t}),$ (23)
which assumes that the variational distribution of ${\mathbf{x}}_{t}$ can be approximated by the corresponding filtering distribution. Additionally, we assume that the following filtering distribution is approximately equal to its one-step backward smoothing counterpart, i.e.,
Assumption 2.
$p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t}),$ (24)
The two assumptions are reasonable and detailed justification can be found in Appendix B.
With these two assumptions, we can significantly streamline the evaluation of the ELBO, which is summarized in the following proposition.
Proposition 1.
Under the two approximations, and all the assumed variational distributions, together with the model joint distribution in Eq. (18), the ELBO can be reformulated as follows:
$\mathcal{L}\approx$ $\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))$ (25)
$-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$
where the first term (log-likelihood) can be analytically evaluated using the EnKF (detailed in Section III-C). The three KL divergence terms can also be computed in closed form, due to the Gaussian nature of the prior and variational distributions [49].
Proof.
The proof can be found in Appendix B. ∎
III-C EnKF-aided ELBO Evaluation
We now demonstrate how to evaluate $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (25) using the EnKF. The general idea can be derived from Bayesian filtering. We first notice that
$p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ $=\int p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\mathrm{d}{\mathbf{x}}_{t},$ (26)
where the prediction distribution, $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})=$
Consequently, we employ the EnKF, a widely used method for handling nonlinearities and high-dimensional state spaces [51], to obtain the filtering distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$ at each time step $t$. Further details are provided below.
@@ III-C1 Prediction
Given the ensemble of states $\{{\mathbf{x}}_{t-1}^{(i)}\}_{i=1}^{N}$ from the posterior distribution $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})$ at time $t-1$, and conditioning on $\mathbf{w}\sim q(\mathbf{w})$ and $\tilde{f}\sim q(\tilde{f})$, we utilize the state transition in Eq. (20) to perform the prediction step. This generates $N$ predicted samples $\{\bar{{\mathbf{x}}}_{t}^{(i)}\}_{i=1}^{N}$ as follows:
$\bar{{\mathbf{x}}}_{t}^{(i)}\sim{\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1}^{(i)})+\bm{\beta}_{t},\ \mathbf{Q}\right),i=1,2,\ldots,N.$ (28)
Therefore, the prediction distribution in Eq. (27) is approximated as:
@@ III-C2 Update
Observations $\mathbf{y}_{t}$ are then assimilated into the predictive ensemble to update the state estimate. Specifically, the predictive ensemble is updated using:
@@
Thus we can obtain the set of $N$ updated samples $\{{\mathbf{x}}_{t}^{(i)}\}_{i=1}^{N}$ from the filtering distribution at time step $t$. These updated samples allow us to recursively compute the samples and prediction distribution for subsequent time steps.
@@
The prediction and update steps are inherently differentiable, as the sampled latent states become differentiable with respect to the state transition function $\bm{f}(\cdot)$ and the noise covariances $\mathbf{Q}$ and $\mathbf{R}$. This property allows for a straightforward gradient-based optimization (e.g. Adam [52] in Algorithm 1), facilitating a practical learning framework for parameter estimation and model tuning [50, 37].
@@
6:         Get prediction samples using Eq. (28)
7:         Get empirical moments $\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t}$ using Eq. (30)
8:         Get Kalman gain: $\bar{\mathbf{G}}_{t}\!=\!\bar{\mathbf{P}}_{t}\bm{C}^{\top}(\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}\!\!+\!\mathbf{R})^{-1}$
9:         Get updated samples using Eq. (31)
10:         Evaluate the log-likelihood using Eq. (33), and
$L_{\ell}=L_{\ell}+\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$
11:     end for
12:     Evaluate $\mathcal{L}$ based on $L_{\ell}$ and Eq. (25)
13:     Maximize $\mathcal{L}$ and update $\bm{\vartheta}$, ${\bm{\zeta}}$ using Adam [52]
14:end while
15:Output: EnKF particles $\{{\mathbf{x}}_{0:T}^{(i)}\}_{i=1}^{N}$, model parameters $\bm{\vartheta}$, and variational parameters ${\bm{\zeta}}$.
Based on the outlined prediction and update steps, the likelihood $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (26) can be easily evaluated. Specifically, at each time step $t$, we have:
Now, we can evaluate our variational lower bound, $\mathcal{L}$, as defined in Eq. (25). At each step, we first utilize the reparameterization trick (see e.g. Eq. (31b)) to sample from the Gaussian variational distributions $q(\mathbf{w})$ and $q(\tilde{f})$. This allows us to obtain an unbiased estimate of the expected log-likelihood, $\mathbb{E}_{q(\mathbf{w},\tilde{f})}[\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})]$ numerically. Due to the reparameterization trick [53], $\mathcal{L}$ is differentiable with respect to the model parameters $\bm{\vartheta}=\{\bm{\theta}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and the variational parameters ${\bm{\zeta}}=\{\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m}_{\mathrm{w}},\bm{\sigma}_{\mathrm{w}}^{2},\mathbf{m},\mathbf{S},{\vec{\mathbf{z}}}\}$. Consequently, we leverage modern differentiation tools, such as PyTorch, to automatically compute the gradients through backpropagation through time (BPTT) and apply gradient-based optimization methods (e.g., Adam) to maximize $\mathcal{L}$ [54, 52]. The detailed routine for implementing our EnKF-aided ETGPSSM is summarized in Algorithm 1.
@@
The proposed ETGPSSM in Algorithm 1 improves upon existing GPSSMs in two primary aspects of efficiency. First, its computational complexity is reduced from $\mathcal{O}(d_{x}M^{3})$ to $\mathcal{O}(M^{3})$ by utilizing a single GP—while the shallow neural network adds minimal computational overhead. Second, the use of a single GP and shallow neural network prevents the quadratic growth of variational parameters, as discussed in Section II-C. Moreover, by avoiding explicit parameterization of the latent states and instead leveraging the EnKF, the proposed method further reduces parametric complexity. Together, these design choices contribute to both parameterization and computational efficiency. Further empirical validation of these improvements is provided in Section IV-B.
In the next section, we empirically evaluate the performance of ETGPSSM across various datasets.
IV Experiments
We first demonstrate the advantages of ETGPSSM in learning non-stationary dynamical systems in Section IV-A. Next, in Section IV-B, we highlight the computational and parametric efficiency and filtering performance of ETGPSSM in chaotic high-dimensional dynamical systems. Finally, Section IV-C showcases the superiors prediction performance of ETGPSSM using multiple real-world time series datasets. More experimental details can be found in the accompanying source code, which is publicly available online. [^4] https://github.com/zhidilin/gpssmProj/tree/main/high_dim_GPSSM
IV-A Non-Stationary Dynamical System Learning
@@
This subsection demonstrates the superior modeling capabilities of ETGPSSM over existing GPSSM and neural network-based methods by evaluating the learning performance on the following underlying dynamical system:
@@
with a slope function:
@@
The slope and oscillatory functions create non-stationary pattern across different regions of the “kink” function (see Fig. 4).
Using this underlying SSM, we generate training data ${\mathbf{y}}_{1:T}$ by fixing $\sigma_{\mathrm{Q}}^{2}$ at 0.05 and varying $\sigma_{\mathrm{R}}^{2}$ systematically across the set $\{0.0008,0.008,0.08,0.8\}$, resulting in four sets of $T=600$ observations each for training purposes. To demonstrate the advantages of ETGPSSM, we compare it with the state-of-the-art (SOTA) neural network-based AD-EnKFs [50] and GP-based EnVI [37]. Consistent with previous works [31, 37], we retain the true emission model and allow the transition model to be learned in all methods, ensuring a fair comparison within the latent space. The system dynamics learning results across different methods are shown in Fig. 4, where our method is represented in the last two columns. The suffixes “dnn” and “bnn” in brackets indicate the network types—Bayesian neural network (BNN) or deep neural network (DNN)—used in our model (see Fig. 3).
The results in Fig. 4 demonstrate that ETGPSSMs consistently achieve superior system dynamics learning performance. Under small observational noise levels ($\sigma_{\mathrm{R}}^{2}=\{0.0008,0.008\}$), ETGPSSMs—particularly ETGPSSM with DNN—excel in learning the dynamics of non-stationary systems. In contrast, their competitor, AD-EnKFs, fails to capture the smooth left segment of the kink function and performs well only on the more zigzag right segment. Meanwhile, EnVI, the SOTA GPSSM, fails to capture the non-stationary behavior of the system and is unable to model the zigzag region effectively. These results highlight the strength of ETGPSSMs in integrating the representational flexibility of neural networks with the GPs, making them particularly well-suited for modeling non-stationary dynamics.
When the observational noise interference becomes more significant $(\sigma_{\mathrm{R}}^{2}=\{0.08,0.8\})$, the learning performance of all methods deteriorates. This is due to the higher estimation errors in the latent states, which hinder accurate learning of the system dynamics. Nevertheless, ETGPSSMs continue to demonstrate superior learning performance. Notably, in this scenario, BNN-based methods, particularly ETGPSSM (bnn), exhibit superior performance and more reliable uncertainty quantification compared to their DNN-based counterparts. These results highlight not only the effectiveness of ETGPSSM’s non-stationary modeling capabilities but also the regularization benefits inherent in Bayesian methods, which contribute to enhanced robustness against significant observational noise.
@@ IV-B High-Dimensional Latent State Estimation
In this subsection, we focus on latent state estimation in the Lorenz-96 system, a mathematical model for studying chaotic behavior in weather and climate dynamics [55]. The system consists of $d_{x}\in\mathbb{Z}_{+}$ coupled ordinary differential equations (ODEs):
@@
Here the constant forcing parameter $F$ is set to $8$, resulting a fully chaotic dynamic [55], where small variations in initial conditions lead to vastly different trajectories. The dynamics model is discretized using the Euler method, and at each time step $t$, observations are generated through a linear emission model ($\bm{C}\!=\!\bm{I}$) with Gaussian additive noise, where the noise level is significant with the covariance given by $\mathbf{R}=4\bm{I}$.
We first empirically demonstrate the efficiency of ETGPSSM in high-dimensional latent state spaces, as illustrated in Fig. 5. This figure highlights both the computational efficiency and the parameter scalability of ETGPSSM in comparison to conventional GPSSMs across varying latent dimensions. The shallow neural network used in ETGPSSM is a three-layer fully connected architecture with ReLU activation and hidden dimensions of $128$ and $64$, which maps from $\mathbb{R}^{d_{x}}$ to the flow parameter space $\Theta\!\subseteq\!\mathbb{R}^{2d_{x}}$. As shown in the left panel, the number of parameters in ETGPSSM increases linearly with the latent state dimension $d_{x}$, in contrast to the quadratic parameter growth observed in GPSSMs. This linear scalability is achieved through the use of a single shared GP combined with normalizing flows, which drastically reduces the total number of parameters compared to GPSSMs that rely on independent GPs for each state dimension.
@@
The right panel of Fig. 5 illustrates the computational efficiency of ETGPSSM by comparing the one-sample execution time for transition evaluation (i.e., Eq. (1a)). Notably, ETGPSSM maintains consistently low running times across increasing state dimensions, whereas GPSSM exhibits rapid growth in computational cost. For instance, at $d_{x}=100$, ETGPSSM completes the evaluation in 0.36 seconds, significantly outperforming GPSSM, which requires 19.56 seconds. This efficiency stems from the reduced computational complexity of ETGPSSM, which scales as $\mathcal{O}(M^{3})$ due to its use of a single GP, in contrast to GPSSM’s $\mathcal{O}(d_{x}M^{3})$ complexity. These results demonstrate that ETGPSSM achieves more scalable parameter growth with increasing state dimensionality while maintaining computational tractability in high-dimensional settings, making it a practical choice for modeling high-dimensional chaotic systems like the Lorenz-96 system.
@@
As illustrated in Fig. 6, ETGPSSM achieves a root mean square error (RMSE) of 1.2028, outperforming AD-EnKF, which yields an RMSE of 1.6187. This result highlights the superior filtering performance of ETGPSSM in high-dimensional chaotic systems. Additionally, the RMSE of the observations (1.9855) serves as a baseline, demonstrating that ETGPSSM significantly reduces the estimation error compared to the raw observations. Although EnKF yields the lowest RMSE of 0.5468, it relies on full model knowledge and does not learn system dynamics. In contrast, ETGPSSM simultaneously learns dynamics and infers latent states, offering a more versatile and effective solution for high-dimensional complex systems.
TABLE I: Prediction performance in terms of RMSE of the different models on the system identification datasets. Mean and standard deviation of the prediction results are shown as an averaged result over ten different seeds. The three lowest RMSE values are highlighted in shades of green, with a deeper one indicating lower RMSE.

### v3

@@ III-C EnKF-aided ELBO Evaluation
Consequently, we employ the EnKF, a widely used method for handling nonlinearities and high-dimensional state spaces [50], to estimate the filtering distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$ at each time step $t$. The details are as follows.
@@ III-C1 Prediction
Given the ensemble of states $\{{\mathbf{x}}_{t-1}^{(i)}\}_{i=1}^{N}$ from the posterior distribution $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})$ at time $t-1$, and conditioning on $\mathbf{w}\sim q(\mathbf{w})$ and $\tilde{f}\sim q(\tilde{f})$, we use the state transition in Eq. (29) to generate $N$ predictive samples:
$\!\!\bar{{\mathbf{x}}}_{t}^{(i)}\!\sim\!{\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1}^{(i)})+\bm{\beta}_{t},\ \mathbf{Q}\right)$ (30)
for $1\leq i\leq N$. The prediction distribution in Eq. (28) is then approximated by the predictive ensemble:
@@ III-C2 Update
Observations $\mathbf{y}_{t}$ is incorporated to update the predictive ensemble. Each sample is updated via:
@@
This yields the updated ensemble $\{{\mathbf{x}}_{t}^{(i)}\}_{i=1}^{N}$ which is used to recursively propagate forward through time.
@@
The prediction and update steps are fully differentiable, as the sampled latent states become differentiable with respect to the state transition function $\bm{f}(\cdot)$ and the noise covariances $\mathbf{Q}$ and $\mathbf{R}$. This allows for end-to-end gradient-based optimization (e.g.using Adam [51], as described in Algorithm 1), making the framework practical for joint learning and inference [52, 37].
Based on the outlined prediction and update steps, the likelihood $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (27) can be easily evaluated. At each time step $t$, we have:
We now proceed to evaluate the variational lower bound $\mathcal{L}$, as defined in Eq. (26). At each iteration, we apply the reparameterization trick (see e.g. Eq. (33b)) to sample from the Gaussian variational distributions $q(\mathbf{w})$ and $q(\tilde{f})$, enabling unbiased Monte Carlo estimates of the expected log-likelihood term $\mathbb{E}_{q(\mathbf{w},\tilde{f})}[\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})]$. Thanks to the reparameterization trick [53], the ELBO $\mathcal{L}$ remains differentiable with respect to both the model parameters $\bm{\vartheta}=\{\bm{\theta}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and the variational parameters ${\bm{\zeta}}=\{\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m}_{\mathrm{w}},\bm{\sigma}_{\mathrm{w}}^{2},\mathbf{m},\mathbf{S},\mathbf{Z}\}$. We leverage modern automatic differentiation tools such as PyTorch to compute gradients via backpropagation through time (BPTT), and apply gradient-based optimization methods (e.g., Adam [51]) to maximize $\mathcal{L}$ [54]. The complete procedure for training ETGPSSM with EnKF-based latent state inference is summarized in Algorithm 1.
@@
6:    EnKF
7:         Get prediction samples using Eq. (30)
8:         Get empirical moments $\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t}$ using Eq. (32)
9:         Get Kalman gain: $\bar{\mathbf{G}}_{t}\!=\!\bar{\mathbf{P}}_{t}\bm{C}^{\top}(\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}\!\!+\!\mathbf{R})^{-1}$
10:         Get updated samples using Eq. (33)
11:         Evaluate the log-likelihood using Eq. (35), and
$L_{\ell}\leftarrow L_{\ell}+\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\vspace{-.1in}$
12:
13:     end for
14:     Evaluate $\mathcal{L}$ based on $L_{\ell}$ and Eq. (26)
15:     Maximize $\mathcal{L}$ and update $\bm{\vartheta}$, ${\bm{\zeta}}$ using Adam [51]
16:end while
17:Output: EnKF particles $\{{\mathbf{x}}_{0:T}^{(i)}\}_{i=1}^{N}$, model parameters $\bm{\vartheta}$, and variational parameters ${\bm{\zeta}}$.
@@
The proposed ETGPSSM in Algorithm 1 improves upon existing GPSSMs in two key aspects of efficiency. First, it reduces the computational complexity from $\mathcal{O}(d_{x}M^{3})$ to $\mathcal{O}(M^{3})$ by relying on a single GP, while the added cost of a shallow neural network remains negligible. Second, by sharing the GP and flow network across dimensions, it avoids the quadratic scaling of variational parameters associated with independent GPs, as discussed in Section II-C. Furthermore, by replacing explicit variational parameterization of the latent states with EnKF-based inference, the method significantly reduces parametric overhead. Together, these design choices lead to substantial gains in both computational and modeling efficiency. Empirical results validating these advantages are presented in Section IV-B.
In the next section, we empirically evaluate the performance of ETGPSSM across a range of dynamical system modeling tasks.
@@
IV Experiments
We begin by evaluating ETGPSSM’s ability to model non-stationary dynamical systems in Section IV-A. We then assess its computational and parametric efficiency, as well as filtering accuracy, in chaotic high-dimensional settings in Section IV-B. Lastly, Section IV-C highlights the model’s predictive performance across a range of real-world time series datasets. Unless otherwise stated, all GPSSMs use the squared exponential (SE) kernel [9], with $20$ inducing points initialized at random. The shallow neural network used in ETGPSSM is a fully connected three-layer architecture with ReLU activations and hidden dimensions $128$ and $64$, mapping from $\mathbb{R}^{d_{x}}$ to the flow parameter space $\Theta\!\subseteq\!\mathbb{R}^{2d_{x}}$. Additional implementation details are available in the publicly released code. [^5] https://github.com/zhidilin/gpssmProj/tree/main/high_dim_GPSSM
IV-A Non-Stationary Dynamical System Learning
This subsection evaluates the modeling capabilities of ETGPSSM on a non-stationary dynamical system, comparing its performance to state-of-the-art GPSSM and neural network-based methods. The underlying system is defined as:
@@ IV-A Non-Stationary Dynamical System Learning
with a slope modulation:
@@
The kink function is a standard benchmark for evaluating GPSSMs [27, 31, 37]. Here, we extend it with non-stationary slope and oscillatory terms to increase modeling complexity (see Fig. 4). We generate training data ${\mathbf{y}}_{1:T}$ by fixing $\sigma_{\mathrm{Q}}^{2}$ at $0.05$ and varying $\sigma_{\mathrm{R}}^{2}$ systematically across the set $\{0.0008,0.008,0.08,0.8\}$, producing four noise regimes. ETGPSSM is compared to two leading baselines: the neural network-based AD-EnKF [52] and the GP-based EnVI [37]. Consistent with previous works [31, 37], all methods share the true emission model while learning the transition function, ensuring a fair comparison. The system dynamics learning results across different methods are shown in Fig. 4, where our method is represented in the last two columns. The suffixes “dnn” and “bnn” in brackets indicate the network types—Bayesian neural network (BNN) or deep neural network (DNN)—used in our model (see Fig. 3).
The results in Fig. 4 show that, across all noise levels, ETGPSSM achieves consistently superior system identification performance. In the low-noise regimes($\sigma_{\mathrm{R}}^{2}=\{0.0008,0.008\}$), the DNN-based ETGPSSM excels at capturing both smooth and oscillatory regions of the transition function. By contrast, AD-EnKF struggles with the smooth left segment and performs well only on the right, highly nonlinear region. EnVI, the strongest GPSSM baseline, fails to capture the non-stationarity and underfits the high-frequency component.
Under high noise $(\sigma_{\mathrm{R}}^{2}=\{0.08,0.8\})$, performance declines across all methods due to increased latent state uncertainty. Nevertheless, ETGPSSM remains robust. Notably, in this scenario, BNN-based methods, particularly ETGPSSM (bnn), outperforms the DNN-based version, benefiting from regularization and better uncertainty quantification. These results emphasize the importance of modeling flexibility and demonstrate the advantage of combining neural flows with GP priors.
@@ IV-B High-Dimensional Latent State Estimation
In this subsection, we evaluate latent state estimation in the Lorenz-96 system, a standard benchmark for studying chaotic dynamics in weather and climate modeling [55]. The system consists of $d_{x}\in\mathbb{Z}_{+}$ coupled ordinary differential equations (ODEs):
@@
where $F=8$ induces fully chaotic behavior, with small differences in initial conditions leading to divergent trajectories [55]. The dynamics model is discretized using the Euler method, and at each time step $t$, observations are generated through a linear emission model ($\bm{C}\!=\!\bm{I}$) with Gaussian additive noise, where the noise level is significant with the covariance given by $\mathbf{R}=4\bm{I}$.
We first assess the scalability of ETGPSSM in high-dimensional settings. Fig. 5 compares its parameter growth and runtime to standard GPSSMs, using $100$ inducing points. As shown in the left panel, ETGPSSM exhibits linear growth in parameter count with increasing $d_{x}$, while GPSSM grows quadratically. This is enabled by ETGPSSM’s use of a shared GP combined with normalizing flows, in contrast to GPSSM’s reliance on independent GPs per dimension.
The right panel of Fig. 5 illustrates the computational efficiency of ETGPSSM by comparing the one-sample execution time for transition evaluation (i.e., Eq. (1a)). Notably, ETGPSSM maintains consistently low running times across increasing state dimensions, whereas GPSSM exhibits rapid growth in computational cost. For instance, at $d_{x}=100$, ETGPSSM completes the evaluation in $0.36$ seconds, significantly outperforming GPSSM, which requires $19.56$ seconds. This efficiency stems from the reduced computational complexity of ETGPSSM, which scales as $\mathcal{O}(M^{3})$ due to its use of a single GP, in contrast to GPSSM’s $\mathcal{O}(d_{x}M^{3})$ complexity. These results demonstrate that ETGPSSM remains computationally and parametrically tractable in high-dimensional latent spaces.
@@
As depicted in Fig. 6, our ETGPSSM achieves an RMSE of $1.2028$, outperforming AD-EnKF, which yields an RMSE of $1.6187$. Additionally, note that the RMSE of the observations (1.9855) serves as a baseline, demonstrating that ETGPSSM significantly reduces the estimation error compared to the raw observations. This result highlights the superior filtering performance of ETGPSSM in high-dimensional chaotic systems. While EnKF achieves the lowest RMSE of $0.5468$, it assumes full knowledge of the dynamics and performs no learning. In contrast, ETGPSSM simultaneously learns the system and infers latent states, offering a more general and powerful solution for complex, high-dimensional systems.
IV-C Real-World Time Series Prediction
In this subsection, we evaluate the prediction performance of ETGPSSM on several real-world time series datasets [^7] https://homes.esat.kuleuven.be/~smc/daisy/daisydata.html commonly used for system identification (see Table I). In every dataset, the first half of the sequence is utilized as training data, with the remaining portion designated for testing. Standardization of all datasets is conducted based on the training sequence. Table I reports the series prediction results, wherein the RMSE is averaged over 50-step ahead forecasting. We compare ETGPSSM against state-of-the-art methods, including neural network-based approaches—DKF [7] and AD-EnKF [52]—as well as various GPSSMs: vGPSSM [29], CO-GPSSM [36], PRSSM [30], ODGPSSM [39], VCDT [31], and EnVI [37]. Following previous works, the latent space dimension is set to $d_{x}=4$ for all the methods.
TABLE I: The 50-step ahead prediction performance in terms of RMSE of different models (all with $d_{x}=4$) on system identification datasets. Mean and standard deviation of the prediction results are shown as an averaged result over ten different seeds. The three lowest RMSE values are highlighted in shades of green, with a deeper one indicating lower RMSE.